OpenStudy (tennistar):

Find the value of 2 times the square root of 3, all over the square root of 10. in simplest form.

OpenStudy (anonymous):

THis is called rationalizing the denominator. \[\frac{2\sqrt{3}}{\sqrt{10}} = \frac{2\sqrt{3}}{\sqrt{10}} \times \frac{\sqrt{10}}{\sqrt{10}} = ?\]

OpenStudy (anonymous):

 Again, you need to simplify the expression by making the denom root-free. Can you go from there?